Question 30 0 out of 6 points Which of the following accurately represents a 95% confidence interval for a sample of 5 individuals whose blood glucose levels were 98, 111, 105, 120, and 147? (78.9, 153.5) (153.5, 78.9) (78.9, 153.5) Selected Answer: Answers: (63.4, 169.0) (169.0, 63.4)Hello everyone, I tried calculating using t and z score and didn't get the answer.
Answer #1

Solution: We are required to find the 95% confidence interval for the population mean.

The 95% confidence interval for the population mean is:

Vn

Where:

\bar{x}=116.2 is the sample mean of given 5 observations

t_{\frac{0.05}{2}}=2.776 is the critical value at 0.05 significance level for df=n-1=5-1=4

s=19.018 is the sample standard deviation of given 5 observations

n=5 is the number of individuals in the given sample

\therefore 116.2\pm 2.776\left ( \frac{19.018}{\sqrt{5}} \right )

116.2\pm 23.6

\left [ 116.2- 23.6,116.2+ 23.6 \right ]

\left [ 92.6,139.8 \right ]

Therefore, the 95% confidence interval for the population mean is \left ( 92.6,139.8 \right )

But none of the given options is what we got.

There is a mistake with the given correct answer \boldsymbol{\left ( 63.4,169.0 \right )} .

Let me explain it:

The answer given above has been calculated using the below formula:

\bar{x}\pm t_{\frac{0.05}{2}} \times s

116.2\pm 2.776 \times 19.018

116.2\pm 52.8

\left ( 116.2- 52.8,116.2+ 52.8 \right )

\left ( 63.4,169.0 \right )

So using the formula \bar{x}\pm t_{\frac{0.05}{2}} \times s , we got exactly what has been shown the correct answer. But the formula used is wrong because they have forgotten to divide standard deviation by \sqrt{n} . That is the reason you were unable to figure out the correct answer.

Hope this helps!